OpenStudy (anonymous):

Change 4x – 2y = 9 into slope-intercept form

OpenStudy (anonymous):

-4x over. -2y=9-4x divide by -2 y=2x-4.5

OpenStudy (anonymous):

medal please? :)

OpenStudy (anonymous):

tnx.

OpenStudy (anonymous):

you got that wrong....

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

you probably want to keep things as fractions

OpenStudy (anonymous):

y=mx+B

jimthompson5910 (jim_thompson5910):

so instead of -4.5, you would write -9/2

OpenStudy (anonymous):

It would be 2y = -4x+9 y=-2x+4.5

OpenStudy (anonymous):

yea, im not wrong though..

OpenStudy (anonymous):

yeh 4.5 is 9/2

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Solve 4 x - 2 y = 9 for y.\[\begin{array}{c} y=2 x-\frac{9}{2} \\ \end{array} \]The slope is 2 and the y intercept is -(9/2)